Suppose that the time, X, that a tourist must wait to board a trolley car in San Francisco is a uniformly distributed random variable over the interval from 0 minutes to 10 minutes.
a.What is the probability that the tourist’s wait-time is exactly 0 minutes?
b.What is the probability that a tourist must wait longer than 8 minutes to board the trolley car?
c.What is the expected wait-time for a tourist?

Answer by stanbon(75887)

Suppose that the time, X, that a tourist must wait to board a trolley car in San Francisco is a uniformly distributed random variable over the interval from 0 minutes to 10 minutes.
a.What is the probability that the tourist’s wait-time is exactly 0 minutes?
P(X=0) = 1/10
--------------
b.What is the probability that a tourist must wait longer than 8 minutes to board the trolley car?
P(X = 9 or 10) = 2/10 = 1/5
---------------
c.What is the expected wait-time for a tourist?
E(X) = (1/10)(1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + 9 + 10)
E(x) = (1/10)[(10/2)(11) = 5.5
===================
Cheers,
Stan H.